# Thermal and non-thermal dark matters with gravitational neutrino reheating

### v1

@@ II.1 The framework
We show that it is the weak decay width, $\Gamma_{3}\simeq\beta^{2}M_{3}/(8\pi)\ll\Gamma_{1},\Gamma_{2}$ of $\nu_{R}^{3}$, with $\beta^{2}\simeq(m_{1}M_{3}/v^{2})$ compared to the other two RHNs (see Appendix-A), that plays crucial role for successful reheating to occur in present scenario. Since $\beta^{2}\propto m_{1}$, the non-vanishing $\Gamma_{3}$ immediately sets the mass of the lightest active neutrino mass $m_{1}$. In other words, to have a neutrino to control the reheating process along with the successful baryogenesis, we indeed require the lightest active neutrino mass to be non-vanishing. This shows a clear correlation between the neutrino parameter and the reheating parameter.
@@ II.4.2 Neutrino heating : $\mathbf{\beta}^{c}_{\nu}\leq\beta\leq\beta^{c}_{\phi}$
Considering details of the different reheating histories discussed above, we now depicts the parameter regions of $\{\beta,M_{3},M_{1},w_{\phi},T_{\rm re}\}$ where different histories are realized. In the bottom panel of Fig.3, we have plotted $T_{\rm re}$ as a function of $\beta$ within the allowed range of $M_{3}$. The different color lines correspond to different values of $M_{3}$. The corresponding mass range is taken within ($10^{10},2\times 10^{13}$) GeV for $w_{\rm\phi}=1/2$ and, ($5\times 10^{10},2\times 10^{13}$) GeV for $w_{\rm\phi}=9/11$. From Fig.3, for a given $\beta$, $T_{\rm re}$ increases with increasing mass of $\nu^{3}_{\rm R}$ $M_{3}$, due to its increase of entropy injection. However, it worth remembering that the gravitational production rate of neutrino is $\propto\left(1-\frac{4M^{2}_{3}}{k^{2}\gamma^{2}m^{2}_{\rm\phi}}\right)^{3/2}$. As a result, for $M_{3}\approx\gamma\,m^{\rm end}_{\rm\phi}\simeq 2\times 10^{13}$ GeV kinetic suppression leads to negligible neutrino production, and consequently reheating would not be possible due to significant kinematic suppression (depicted by red shaded region). On the other hand, the lowest possible mass $M_{3}$ can be determined from the condition $T_{\rm re}(\beta^{\rm c}_{\rm\nu})\sim T_{\rm re}^{\rm min}$, and the final expression of $M^{\rm min}_{3}$ is
@@ III Constraints from $\Delta N_{\rm eff}$ considering PGWs
PGWs is one of the profound predictions of inflationary paradigm Grishchuk:1974ny ; Guzzetti:2016mkm ; Starobinsky:1979ty . Due to extremely weak coupling, it can pass through the universe almost unhindered and hence acts as a unique probe of the very early universe such as reheating phase Mishra:2021wkm ; Benetti:2021uea ; Haque:2021dha ; Maity:2024cpq . The PGW amplitude and its spectral evolution are expected to be sensitive to the energy scale of the inflation and the post-inflationary phases. Particularly during the reheating phase, it has been observed that the modes between $k_{\rm re}<k<k_{\rm end}$ which re-enter the horizon during reheating, PGW spectrum shows a tilt depending on the background equation of state. Once we fixed the inflationary energy scale and $w_{\phi}$, $T_{\rm re}$ regulates the duration of the tilted spectrum. The tilted spectrum’s duration extends as we reduce $T_{\rm re}$. Nonetheless, if one considers the extra relativistic degrees of freedom arising from the GWs, the lesser value of $T_{\rm re}$ for a stiff equation can be constrained. We are going to discuss how small $T_{\rm re}$ after we fix the EoS and the inflation’s energy scale, we can consider being consistent with the $\Delta N_{\rm eff}$. The additional relativistic degrees of freedom at the moment of BBN or CMB decoupling are denoted by $\Delta N_{\rm eff}$. The expression for $\Delta N_{\rm{eff}}$ in the case of the GWs takes the following form Jinno:2012xb
@@
where, $\mu(w_{\phi})=(1+3\omega_{\phi})^{\frac{4}{1+3\,w_{\phi}}}\,\Gamma^{2}\left(\frac{5+3\,w_{\phi}}{2+6\,w_{\phi}}\right)$, which typically assumes ${\cal O}(1)$ value. $k_{\rm end}$ is the largest frequency mode that reenters the horizon at the end of the inflation, which typically contributes to the $\Delta N_{\rm{eff}}$. Note that if there is no intermediate neutrino (matter) domination such as the case of usual GRe, the form of the above spectrum Eq.39 would be the same with $k_{\nu}$ replaced by $k_{\rm re}$. This essentially suggests that due to late matter domination, spectrum in Eq.39 acquires a large suppression factor $({k_{\rm re}}/{k_{\nu}})^{\frac{12\,w_{\phi}}{(1+3\,w_{\phi})}}<<1$ for high frequency modes. For example taking $w_{\rm\phi}=9/11$, we obtain $T_{\rm re}^{\rm GRe}\sim 4\times 10^{3}$ GeV and $k_{\rm re}\sim 10^{-3}$ Hz. Whereas taking $M_{3}=5\times 10^{13}$ GeV and keeping the same $w_{\rm\phi}\,,T_{\rm re}$ as above, $\nu$GRe yields $k_{\rm\nu}\sim 10^{-2}$ Hz. Utilizing those values we obtain suppression factor to be $\sim 10^{-3}$. Therefore, neutrino dominating case always abide by the $\Delta N_{\rm eff}$ constraint for all equation state $w_{\rm\phi}>1/3$.
@@ V Leptogenesis and constraints
RHNs are gravitationally produced from the inflaton. Therefore, all of them undergo CP-violating out-of-equilibrium decay and produce lepton asymmetry. By the well-known non-perturbative sphaleron processes, those lepton asymmetries are then converted into the baryon asymmetry. For our analysis we considered the following mass hierarchy $M_{1}\lesssim m^{\rm end}_{\rm\phi}\ll M_{2}$. Note that due to its very small decay width, $\nu_{\rm R}^{3}$ does not contribute significant asymmetry. The lepton asymmetry will, therefore, predominantly be produced by the decays of $\nu^{1}_{\rm R}$. The CP asymmetry parameter $(\epsilon_{\Delta L})$ generated from the decay of $\nu^{1}_{\rm R}$, using the seesaw mechanism Buchmuller:2004nz ; Barman:2022qgt ; Kaneta:2019yjn ; Barman:2023opy is expressed as ,
@@
$\simeq\frac{3\,\delta_{\rm eff}}{16\,\pi}\frac{m_{3}\,M_{1}}{v^{2}}\,$
@@
Where, $\delta_{\rm eff}$ is the effective CP-violating phase in the neutrino mass matrix, which typically assumes the value within $0\leq\delta_{\rm eff}\leq 1$, and we have taken $\delta_{\rm eff}=1$. We take the normal hierarchy of active neutrino mass with $m_{3}=0.05$ eV, and the Higgs vacuum expectation value $v=174$ GeV. As stated earlier, the above lepton asymmetry is converted into the baryon asymmetry $(Y_{\rm B})$ via the electroweak sphaleron processes sy ; ja ,
@@ VI Particle Dark matter phenomenology
In the recent past, the effect of non-standard cosmology has gained significant interest in the DM phenomenological studies Maity:2018dgy ; Garcia:2020eof ; Garcia:2020wiy ; Garcia:2021gsy ; Maity:2018exj ; Haque:2020zco ; Giudice:2000ex ; Barman:2022tzk ; Bhattiprolu:2022sdd ; Harigaya:2014waa ; Harigaya:2019tzu ; Okada:2021uqk ; Ghosh:2022fws . In this paper, we aim to discuss how the present neutrino reheating scenario affects the DM production and constrain its parameter space. We consider both thermal and non-thermal DM production scenarios. Let us consider a real standard model singlet DM $S$ Silveira:1985rk ; Burgess:2000yq ; McDonald:1993ex ; Andreas:2008xy ; Davoudiasl:2004be ; Vangsnes:2021rni ; Lebedev:2021xey ; Guo:2010hq ; Yaguna:2011ei ; Cline:2013gha , which is odd and all the standard model particles are even under the $\mathbb{Z}_{2}$ symmetry. The DM interacts with the SM via the Higgs portal interaction and also with inflaton through universal gravitational interaction. The interaction Lagrangian for DM is,
@@
$\mu_{\rm N}=m_{\rm N}\,m_{\rm s}/(m_{\rm N}+m_{\rm s})$ is the reduced mass of the DM-nucleon system with nucleon mass $m_{\rm N}=1\,$GeV and $f_{\rm N}=0.30$. Taking $\sigma_{\rm SI}$ from various direct-detection experiments such as DarkSide-50 (DS-50) DarkSide-50:2023fcw , XENONnT XENON:2023cxc , LUX-ZEPLIN (LZ) LZ:2022lsv , XLZD Aalbers:2022dzr and we have projected this on ($\lambda_{\rm hs}\,,m_{\rm s}$) plane. As indeed, one can see, some parts of the parameter space are already ruled out by the existing experimental bounds from DS-50 (purple-shaded), XENONnT (magenta-shaded), and LZ (brown-shaded). Reheating, however, provides wider allowed parameter space for WIMP-like dark matter, which may be possible to detect in the future in experiments like XLZD within the gray region. For the FIMP-like DM, however, due to extremely weak coupling, all the parameter spaces are allowed.
@@ VII QCD Axion as a DM
with critical energy density $\rho_{c}=1.05\times 10^{-5}\,h^{2}\,\rm{GeV/cm^{3}}$,$\,s(T_{0})\simeq 2.69\times 10^{3}\,\rm cm^{-3}$ is the entropy density at today and $m_{\rm\mathfrak{a}}^{\rm qcd}\simeq 4.8\times 10^{-11}$ eV.
We consider the situation where axion oscillation begins during the reheating phase, i.e., $T_{\rm osc}>T_{\rm re}$. For such a case, the axion energy density at the present epoch is prescribed as Xu:2023lxw ; Barman:2023icn ; Grin:2007yg ; Visinelli:2009kt ; Arias:2021rer ; Bernal:2021yyb
@@ VIII Conclusions
RM would like to thank the Ministry of Human Resource Development, Government of India (GoI), for financial assistance. MRH wishes to acknowledge support from the Science and Engineering Research Board (SERB), Government of India (GoI), for the SERB National Post-Doctoral fellowship, File Number: PDF/2022/002988. DM wishes to acknowledge support from the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (GoI), through the Core Research Grant CRG/2020/003664. RM would like to thank Nayan Das and Disha Bandyopadhyay for the helpful discussions. DM and RM also thank the Gravity and High Energy Physics groups at IIT Guwahati for illuminating discussions.

### v2

# Thermal and nonthermal dark matters with gravitational neutrino reheating

@@ II.1 The framework
We show that it is the weak decay width, $\Gamma_{3}\simeq\beta^{2}M_{3}/(8\pi)\ll\Gamma_{1},\Gamma_{2}$ of $\nu_{R}^{3}$, with $\beta^{2}\simeq(m_{1}\,M_{3}/v^{2})$ compared to the other two RHNs (see Appendix-A), that plays crucial role for successful reheating to occur in present scenario. Since $\beta^{2}\propto m_{1}$ [^1] Since we use the normal hierarchy of active neutrino masses here, $\beta^{2}\propto m_{1}$. However, $\nu$GRe is also possible if one considers the inverted hierarchy, in which case $\beta^{2}\propto m_{3}$. See Appendix-A, where we construct the Yukawa coupling using both hierarchies., the non-vanishing $\Gamma_{3}$ immediately sets the mass of the lightest active neutrino mass $m_{1}$. In other words, to have a neutrino to control the reheating process along with the successful baryogenesis, we indeed require the lightest active neutrino mass to be non-vanishing. This shows a clear correlation between the neutrino parameter and the reheating parameter.
@@ II.4.2 Neutrino heating : $\mathbf{\beta}^{c}_{\nu}\leq\beta\leq\beta^{c}_{\phi}$
Considering details of the different reheating histories discussed above, we now depict the parameter regions of $\{\beta\,,M_{3}\,,w_{\phi}\,,T_{\rm re}\}$ where different histories are realized. In the bottom panel of Fig.3, we have plotted $T_{\rm re}$ as a function of $\beta$ within the allowed range of $M_{3}$. The different color lines correspond to different values of $M_{3}$. The corresponding mass range is taken within ($10^{10},2\times 10^{13}$) GeV for $w_{\rm\phi}=1/2$ and, ($5\times 10^{10},2\times 10^{13}$) GeV for $w_{\rm\phi}=9/11$. From Fig.3, for a given $\beta$, $T_{\rm re}$ increases with increasing mass of $\nu^{3}_{\rm R}$ $M_{3}$, due to its increase of entropy injection. However, it worth remembering that the gravitational production rate of neutrino is $\propto\left(1-\frac{4M^{2}_{3}}{k^{2}\gamma^{2}m^{2}_{\rm\phi}}\right)^{3/2}$. As a result, for $M_{3}\approx\gamma\,m^{\rm end}_{\rm\phi}\simeq 2\times 10^{13}$ GeV kinetic suppression leads to negligible neutrino production, and consequently reheating would not be possible due to significant kinematic suppression (depicted by red shaded region). On the other hand, the lowest possible mass $M_{3}$ can be determined from the condition $T_{\rm re}(\beta^{\rm c}_{\rm\nu})\sim T_{\rm re}^{\rm min}$, and the final expression of $M^{\rm min}_{3}$ is
@@ III Constraints from $\Delta N_{\rm eff}$ considering PGWs
PGWs is one of the profound predictions of inflationary paradigm Grishchuk:1974ny ; Guzzetti:2016mkm ; Starobinsky:1979ty . Due to extremely weak coupling, it can pass through the universe almost unhindered and hence acts as a unique probe of the very early universe such as reheating phase Mishra:2021wkm ; Benetti:2021uea ; Haque:2021dha ; Maity:2024cpq ; Vagnozzi:2023lwo ; Vagnozzi:2020gtf . The PGW amplitude and its spectral evolution are expected to be sensitive to the energy scale of the inflation and the post-inflationary phases. Particularly during the reheating phase, it has been observed that the modes between $k_{\rm re}<k<k_{\rm end}$ which re-enter the horizon during reheating, PGW spectrum shows a tilt depending on the background equation of state. Once we fixed the inflationary energy scale and $w_{\phi}$, $T_{\rm re}$ regulates the duration of the tilted spectrum. The tilted spectrum’s duration extends as we reduce $T_{\rm re}$. Nonetheless, if one considers the extra relativistic degrees of freedom arising from the GWs, the lesser value of $T_{\rm re}$ for a stiff equation can be constrained. We are going to discuss how small $T_{\rm re}$, after we fix the EoS and the inflation’s energy scale, we can consider being consistent with the $\Delta N_{\rm eff}$. The additional relativistic degrees of freedom at the moment of BBN or CMB decoupling are denoted by $\Delta N_{\rm eff}$. The expression for $\Delta N_{\rm{eff}}$ in the case of the GWs takes the following form Jinno:2012xb
@@
where, $\mu(w_{\phi})=(1+3\omega_{\phi})^{\frac{4}{1+3\,w_{\phi}}}\,\Gamma^{2}\left(\frac{5+3\,w_{\phi}}{2+6\,w_{\phi}}\right)$, which typically assumes ${\cal O}(1)$ value. $k_{\rm end}$ is the largest frequency mode that reenters the horizon at the end of the inflation, which typically contributes to the $\Delta N_{\rm{eff}}$. Note that if there is no intermediate neutrino (matter) domination such as the case of usual GRe, the form of the above spectrum Eq.39 would be the same with $k_{\nu}$ replaced by $k_{\rm re}$. This essentially suggests that due to late matter domination, spectrum in Eq.39 acquires a large suppression factor $({k_{\rm re}}/{k_{\nu}})^{\frac{12\,w_{\phi}}{(1+3\,w_{\phi})}}<<1$ for high frequency modes. For example taking $w_{\rm\phi}=9/11$, we obtain $T_{\rm re}^{\rm GRe}\sim 4\times 10^{3}$ GeV and $k_{\rm re}\sim 10^{-3}$ Hz. Whereas taking $M_{3}=5\times 10^{13}$ GeV and keeping the same $w_{\rm\phi}\,,T_{\rm re}$ as above, $\nu$GRe yields $k_{\rm\nu}\sim 10^{-2}$ Hz. Utilizing those values, we obtain the suppression factor to be $\sim 10^{-3}$. Therefore, the neutrino dominating case always abides by the $\Delta N_{\rm eff}$ constraint for all equation state $w_{\rm\phi}>1/3$.
@@ V Leptogenesis and constraints
RHNs are gravitationally produced from the inflaton. Therefore, all of them undergo CP-violating out-of-equilibrium decay and produce lepton asymmetry. By the well-known non-perturbative sphaleron processes, those lepton asymmetries are then converted into the baryon asymmetry. For our analysis we considered the following mass hierarchy $M_{1}\lesssim m^{\rm end}_{\rm\phi}\ll M_{2}$. Note that due to its very small decay width, $\nu_{\rm R}^{3}$ does not contribute significant asymmetry. The lepton asymmetry will, therefore, predominantly be produced by the decays of $\nu^{1}_{\rm R}$  [^2] Note that the three-body decay such as $\nu^{1,2}_{R}\rightarrow\nu^{3}_{R}+\nu_{L}+\nu_{L}$ is kinematically allowed. However, we checked that the associated production rate of $\nu^{3}_{R}$, $\Gamma_{1\rightarrow 3}\simeq\frac{\beta^{2}\,(y^{\dagger}\,y)_{11/22}}{2\,M_{1/2}}\left(\frac{3}{2}p^{2}+\frac{1}{2}-2\,p-p^{2}\,\ln{p}\right)\,\,,p=M^{2}_{3}/M^{2}_{1/2}$) is very much suppressed compared to the gravitational contributions.. The CP asymmetry parameter $(\epsilon_{\Delta L})$ generated from the decay of $\nu^{1}_{\rm R}$, using the seesaw mechanism Buchmuller:2004nz ; Barman:2022qgt ; Kaneta:2019yjn ; Barman:2023opy is expressed as ,
@@
$\simeq\frac{3\,\delta_{\rm eff}}{16\,\pi}\frac{m_{\rm\nu,max}\,M_{1}}{v^{2}}\,$
@@
Where, $\delta_{\rm eff}$ is the effective CP-violating phase in the neutrino mass matrix, which typically assumes the value within $0\leq\delta_{\rm eff}\leq 1$, and we have taken $\delta_{\rm eff}=1$. We take the normal hierarchy of active neutrino mass with $m_{\rm\nu,max}=m_{3}=0.05$ eV  [^3] For the inverted hierarchy $m_{\rm\nu,max}=m_{2}=0.05$ eV., and the Higgs vacuum expectation value $v=174$ GeV. As stated earlier, the above lepton asymmetry is converted into the baryon asymmetry $(Y_{\rm B})$ via the electroweak sphaleron processes sy ; ja ,
@@ VI Particle Dark matter phenomenology
In the recent past, the effect of non-standard cosmology has gained significant interest in the DM phenomenological studies Maity:2018dgy ; Garcia:2020eof ; Garcia:2020wiy ; Garcia:2021gsy ; Maity:2018exj ; Haque:2020zco ; Giudice:2000ex ; Barman:2022tzk ; Bhattiprolu:2022sdd ; Harigaya:2014waa ; Harigaya:2019tzu ; Okada:2021uqk ; Ghosh:2022fws ; Haque:2021mab ; Ahmed:2022tfm ; Bernal:2022wck ; Bernal:2023ura ; Chowdhury:2023jft . In this paper, we aim to discuss how the present neutrino reheating scenario affects the DM production and constrain its parameter space. We consider both thermal and non-thermal DM production scenarios. Let us consider a real standard model singlet DM $S$ Silveira:1985rk ; Burgess:2000yq ; McDonald:1993ex ; Andreas:2008xy ; Davoudiasl:2004be ; Vangsnes:2021rni ; Lebedev:2021xey ; Guo:2010hq ; Yaguna:2011ei ; Cline:2013gha , which is odd and all the standard model particles are even under the $\mathbb{Z}_{2}$ symmetry. The DM interacts with the SM via the Higgs portal interaction and also with inflaton through universal gravitational interaction. The interaction Lagrangian for DM is,
@@
$\mu_{\rm N}=m_{\rm N}\,m_{\rm s}/(m_{\rm N}+m_{\rm s})$ is the reduced mass of the DM-nucleon system with nucleon mass $m_{\rm N}=1\,$GeV and $f_{\rm N}=0.30$. Taking $\sigma_{\rm SI}$ from various direct-detection experiments such as DarkSide-50 (DS-50) DarkSide-50:2023fcw , XENONnT XENON:2023cxc , LUX-ZEPLIN (LZ) LZ:2022lsv , XLZD Aalbers:2022dzr and we have projected this on ($\lambda_{\rm hs}\,,m_{\rm s}$) plane. As indeed, one can see, some parts of the parameter space are already ruled out by the existing experimental bounds from DS-50 (purple-shaded), XENONnT (magenta-shaded), and LZ (brown-shaded). Reheating, however, provides wider allowed parameter space for WIMP-like dark matter, which may be possible to detect in the future in experiments like XLZD within the gray region. For the FIMP-like DM, however, due to extremely weak coupling, all the parameter spaces are allowed. Note that the indirect DM detection constraint in Higgs portal models is superseded by those from direct detection experiments and the LHC. Therefore, we do not consider them in our analysis.
@@ VII QCD Axion as a DM
with critical energy density $\rho_{c}=1.05\times 10^{-5}\,h^{2}\,\rm{GeV/cm^{3}}$,$\,s(T_{0})\simeq 2.69\times 10^{3}\,\rm cm^{-3}$ is the entropy density at today and $m_{\rm\mathfrak{a}}^{\rm qcd}\simeq 4.8\times 10^{-11}$ eV. To match the observed DM abundance, the initial misalignment angle $\theta_{\rm i}$ could be tuned. Assuming an initial misalignment angle $\theta_{\rm i}\simeq 1.0$, the above equation indicates that an axion mass around $m_{\rm\mathfrak{a}}\sim\mathcal{O}(10^{-5})$ eV and the decay constant $f_{\mathfrak{a}}\sim\mathcal{O}(10^{12})$ GeV is required to match the observed relic abundance.
The initial misalignment angle is deeply connected to the cosmological history of the axion. If the Peccei-Quinn (PQ) symmetry is broken before or during inflation (and not restored afterward), the initial angle $\theta_{i}$ will be uniform across the observable universe, and the angle is randomly chosen from the range $[-\pi,\pi]$. For $\theta_{\rm i}\sim\mathcal{O}(1)$, $m_{\rm\mathfrak{a}}$ should be in the range of $\mu$eV to avoid overproducing the DM abundance. Smaller axion masses (or equivalently, larger values of the decay constant $f_{\mathfrak{a}}$) become viable if small-tuned values of $\theta_{i}$ are considered. This is often referred to as the anthropic axion window Wantz:2009it ; Hertzberg:2008wr . In this pre-inflationary scenario, large values of $f_{\mathfrak{a}}$ are constrained by isocurvature perturbations that impose $f_{\mathfrak{a}}\lesssim 10^{16}$ GeV Arias:2021rer . On the other hand, if the PQ symmetry breaks after inflation, i.e, in a post-inflationary scenario, the initial value of $\theta_{i}$ may be different at different points in space, since we are interested in mode with zero momentum, therefore, we can safely use Eq. 71. When it breaks after the inflationary epoch, topological defects such as strings and domain walls could potentially contribute to axion production, affecting the mass range predictions Arias:2021rer ; Gorghetto:2018ocs ; Kawasaki:2014sqa ; Hiramatsu:2010yu . However, due to the persistent uncertainties regarding the extent of these contributions, our analysis has been limited to axion production through the misalignment mechanism. In this work, we primarily explore the impact of $\nu$GRe on the misalignment mechanism. Depending on the decay scale $f_{a}$, both pre-inflationary ($f_{\mathfrak{a}}\geq H_{\rm end}$) and post-inflationary ($f_{\mathfrak{a}}<H_{\rm end}$) scenarios can emerge.
Now, we consider the situation where axion oscillation begins during the reheating phase, i.e., $T_{\rm osc}>T_{\rm re}$. For such a case, the axion energy density at the present epoch is prescribed as Xu:2023lxw ; Barman:2023icn ; Grin:2007yg ; Visinelli:2009kt ; Arias:2021rer ; Bernal:2021yyb
@@ VIII Conclusions
RM would like to thank the Ministry of Human Resource Development, Government of India (GoI), for financial assistance. MRH wishes to acknowledge support from the Science and Engineering Research Board (SERB), Government of India (GoI), for the SERB National Post-Doctoral fellowship, File Number: PDF/2022/002988. DM wishes to acknowledge support from the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (GoI), through the Core Research Grant CRG/2020/003664. RM would like to thank Nayan Das and Disha Bandyopadhyay for the helpful discussions. DM and RM also thank the Gravity and High Energy Physics groups at IIT Guwahati for illuminating discussions. We thank the anonymous referee for his critical review and comments which helped our paper greatly.